OpenStudy (anonymous):

The defining equation for the function k is k(x) = 3x / (x+1) . Find k(5-h) .

OpenStudy (anonymous):

(9+h)/(h+9) (15-3h)/(-h+6) (3x+9+h) / (x+1) Not enough information is given to answer the question

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

just put x=5-h

OpenStudy (gorv):

@aprilann143 got it ???

OpenStudy (anonymous):

i dont understand how to do that. it gets confusing

OpenStudy (gorv):

k(x) = 3x / (x+1) k(5-h) it suggest tha x is replaced by 5-h

OpenStudy (gorv):

k(5-h) =3*(5-h)/(5-h+1) =(15-3h)/(6-h)

OpenStudy (anonymous):

ty